A compound has been found to be made up of 5.200g chromium, 4.815g sulfur and 9.600g oxygen. Find the empirical formula of this compound.
1
Expert's answer
2020-05-04T01:52:47-0400

Moles of chromium = "\frac{5.2}{52}= 0.1"

Moles of sulfur = "\frac{4.815}{32}=0.15"

Moles of oxygen = "\frac{9.6}{16}= 0.6"

So

"Cr:S:O= 0.1:0.15:0.6"

Multiply by 2

We get "Cr:S:O= 2:3:12"

So empirical formula is -

"Cr_2S_3O_{12}"